OpenStudy (anonymous):

Here's a conceptual question I'd like to clear up. In classical electrodynamics, "waves" lie in Hilbert space, not in real space, right?

OpenStudy (jamesj):

Let's see. Aren't they in C^3 x R (which is not a Hilbert space)

OpenStudy (anonymous):

I'm not sure I understand. Could you clarify?

OpenStudy (jamesj):

I mean that we write the space components x, y, z are complex numbers and time t as a real number. Hence the 4-vector (x,y,z,t) is a member of \[ \mathbb{C}^3 \times \mathbb{R} \]

OpenStudy (anonymous):

Well, I got this part. I'm not sure I understand how this isn't a Hilbert space, though.

OpenStudy (jamesj):

By definition, a Hilbert space is a complete, COMPLEX inner product space. As the time component is not complex, this vector space isn't a Hilbert space. But it does have many of the properties of a Hilbert space. However, I'm not sure thinking of it in terms of Hilbert formalism adds much to the situation.

OpenStudy (anonymous):

Oh, right. Complex. XD Thanks.